Joe mixes 10 ounces of metal that contains 30% gold with 20 ounces of a metal that contains 15% gold. What is the percent of gold in the resulting metal?
F. 30%
G. 22.5%
H. 20%
J. 15%

Answer by alicealc(293)

10 * 30% + 20 * 15% = (10+20) * x%
(omit the % sign for all)
300 + 300 = 30x
600 = 30x
x = 600/30 = 20

so the resulting metal will contain 20% of gold